Fishyr:

Hey can someone help me with this math equation? It's on my Unit Test. If 2a − 1 = 4y, y ≠ 0 where a ≠ 1, what is y in terms of a? 1.y = 2a − 2 2.y = 2a − 4 3.y = 2a − 1a 4.y = yaa + 1

Bounty:

u have to substitute

Fishyr:

thanks

Fishyr:

what next?...

toga:

To solve for y in terms of a, we can rearrange the equation 2a - 1 = 4y to solve for y: 2a - 1 = 4y 2a = 4y + 1 y = (2a - 1) / 4 So, the correct answer is: y = (2a - 1) / 4

Fishyr:

okay, thank you.

toga:

@fishyr wrote:
okay, thank you.
no problem